OpenStudy (anonymous):

What is the distance between the points (–8, –2) and (–7, –2)? A. 1 B. 6 C. 10 D. 15

Mehek (mehek14):

\(\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\) so \(\sqrt{(-8 - (-7))^2+(-2-(-2)^2}\)

Mehek (mehek14):

\(\sqrt{(-1)^2+(0)^2}\)